# Galaxy mass profiles with convolutional neural networks

### v1

@@ Acknowledgements.
JSA thanks the Spanish Ministry of Economy and Competitiveness (MINECO) for support through a grant P/301404 from the Severo Ochoa project CEX2019-000920-S. CB is supported by the Spanish Ministry of Science and Innovation (MICIU/FEDER) through research grant PID2021-122603NBC22. ADC is supported by the Agencia Estatal de Investigación, under the 2023 call for Ayudas para Incentivar la Consolidación Investigadora, grant number CNS2023- 144669 , proyecto ”TINY”. This research made use of computing time available on high performance computing systems at the Instituto de Astrofisica de Canarias. The freely available software pynbody (Pontzen et al., 2013) has been used for part of this analysis. We thank the PIs of the AURIGA and FIRE-2 simulation projects for making their data publicly available at https://wwwmpa.mpa-garching.mpg.de/auriga/data.html and http://flathub.flatironinstitute.org/fire, respectively.References
- Abadi et al. (2015) Abadi, M., Agarwal, A., Barham, P., et al. 2015, TensorFlow: Large-Scale Machine Learning on Heterogeneous Systems, software available from tensorflow.org
- Alsing et al. (2019) Alsing, J., Charnock, T., Feeney, S., & Wandelt, B. 2019, Monthly Notices of the Royal Astronomical Society
- Alsing et al. (2018) Alsing, J., Wandelt, B., & Feeney, S. 2018, Monthly Notices of the Royal Astronomical Society, 477, 2874–2885
@@
- Blank et al. (2019) Blank, M., Macciò, A. V., Dutton, A. A., & Obreja, A. 2019, Monthly Notices of the Royal Astronomical Society, 487, 5476–5489
@@
- Buck et al. (2019) Buck, T., Obreja, A., Macciò, A. V., et al. 2019, Monthly Notices of the Royal Astronomical Society, 491, 3461
@@
- de los Rios et al. (2023) de los Rios, M., Petač, M., Zaldivar, B., et al. 2023, Monthly Notices of the Royal Astronomical Society, 525, 6015
@@
- Dutton et al. (2020) Dutton, A. A., Buck, T., Macciò, A. V., et al. 2020, Monthly Notices of the Royal Astronomical Society, 499, 2648
@@
- Grand et al. (2024) Grand, R. J. J., Fragkoudi, F., Gómez, F. A., et al. 2024, Overview and public data release of the augmented Auriga Project: cosmological simulations of dwarf and Milky Way-mass galaxies
@@
- Ho et al. (2024b) Ho, M., Bartlett, D. J., Chartier, N., et al. 2024b, The Open Journal of Astrophysics, 7
- Ho et al. (2019) Ho, M., Rau, M. M., Ntampaka, M., et al. 2019, The Astrophysical Journal, 887, 25
- Hopkins et al. (2018a) Hopkins, P. F., Wetzel, A., Kereš, D., et al. 2018a, Monthly Notices of the Royal Astronomical Society, 480, 800–863
- Hopkins et al. (2018b) Hopkins, P. F., Wetzel, A., Kereš, D., et al. 2018b, Monthly Notices of the Royal Astronomical Society, 480, 800–863
@@
- Navarro et al. (1997) Navarro, J. F., Frenk, C. S., & White, S. D. M. 1997, The Astrophysical Journal, 490, 493–508
- Nguyen et al. (2023) Nguyen, T., Mishra-Sharma, S., Williams, R., & Necib, L. 2023, Physical Review D, 107
- Papamakarios et al. (2021) Papamakarios, G., Nalisnick, E., Rezende, D. J., Mohamed, S., & Lakshminarayanan, B. 2021, Journal of Machine Learning Research, 22, 1
@@
- Sawala et al. (2016) Sawala, T., Frenk, C. S., Fattahi, A., et al. 2016, Monthly Notices of the Royal Astronomical Society, 457, 1931
- Scarselli et al. (2009) Scarselli, F., Gori, M., Tsoi, A. C., Hagenbuchner, M., & Monfardini, G. 2009, IEEE Transactions on Neural Networks, 20, 61
@@
- Shen et al. (2010) Shen, S., Wadsley, J., & Stinson, G. 2010, Monthly Notices of the Royal Astronomical Society, 407, 1581–1596
@@
- Wang et al. (2015) Wang, L., Dutton, A. A., Stinson, G. S., et al. 2015, Monthly Notices of the Royal Astronomical Society, 454, 83–94
- Wetzel et al. (2023) Wetzel, A., Hayward, C. C., Sanderson, R. E., et al. 2023, The Astrophysical Journal Supplement Series, 265, 44
@@ Appendix A Bivariate Kernel Density Estimation
Kernel density estimation (KDE) is a non-parametric technique used to estimate the PDF of a random variable. Bivariate KDE extends this concept to two dimensions, estimating the joint PDF of two random variables.
@@
Where …, conform a sample of size n of the random variable(s) whose PDF we aim to estimate. In the case of bivariate KDE, represents two values , which are joint samples of the two random variables for which we want to estimate the PDF. is a kernel function and is a bandwidth matrix whose dimensions match those of (in this case, 2x2).
Effectively, the KDE evaluates the PDF at by adding the contributions of each point , weighted by the kernel function , based on the distance between and , scaled by the bandwidth matrix .
We use a 2D Gaussian kernel that follows
where . The bandwidth matrix is calculated by multiplying the covariance matrix of the data by a factor , where n is the size of the sample.
@@ Appendix B MAF Posterior Corner Plots
Appendix C Additional Posterior Calibration Checks
@@
We conduct marginal percentile coverage tests on the model trained on NIHAO and AURIGA simulations simultaneously, assessing the calibration of the confidence intervals specifically for each dimension of the 10-d PDF. We show the tests for the validation set and for FIRE galaxy projections in 12. We find similar results to the TARP, demonstrating that the calibration of the posteriors is correct when marginalizing over any of its dimensions when validating the model on simulations from the suites used to train it. The posterior distributions produced for FIRE galaxies show similar deviations in their calibration in all dimensions.
@@
After checking the calibration of the errors provided by our CNN+MAF model, we now study their magnitudes. In Fig. 13 we plot the magnitude of the uncertainties estimated by our model for the log10 value of the enclosed mass at different radii, for both the training and the validation set of the model trained on NIHAO and AURIGA simulations simultaneously, where we define the uncertainty as half the range between the 16th and 84th percentile of the enclosed mass at each radii according to the predicted posterior. The median uncertainty peaks for both the validation and training set at 0.11 for the mass enclosed within 0.6Rh, and it decreases monotonically with radius, stabilising at 0.078 for masses within radii equal and higher than 1.8Rh. We remind the reader that posterior distributions generated for FIRE galaxies are not correctly calibrated, and therefore the errors shown for this dataset are not representative of the actual deviation between the predicted and the truth value of the enclosed mass.

### v4

@@ Acknowledgements.
JSA thanks the Spanish Ministry of Economy and Competitiveness (MINECO) for support through a grant P/301404 from the Severo Ochoa project CEX2019-000920-S. CB is supported by the Spanish Ministry of Science and Innovation (MICIU/FEDER) through research grant PID2021-122603NBC22. ADC is supported by the Agencia Estatal de Investigación, under the 2023 call for Ayudas para Incentivar la Consolidación Investigadora, grant number CNS2023-144669, proyecto ”TINY”. The authors wish to acknowledge the contribution of the IAC High-Performance Computing support team and hardware facilities to the results of this research. The freely available software pynbody (Pontzen et al., 2013) has been used for part of this analysis. We thank the PIs of the AURIGA and FIRE-2 simulation projects for making their data publicly available at https://wwwmpa.mpa-garching.mpg.de/auriga/data.html and http://flathub.flatironinstitute.org/fire, respectively.References
- Alsing et al. (2019) Alsing, J., Charnock, T., Feeney, S., & Wandelt, B. 2019, MNRAS
- Alsing et al. (2018) Alsing, J., Wandelt, B., & Feeney, S. 2018, MNRAS, 477, 2874–2885
@@
- Arora et al. (2023) Arora, N., Courteau, S., Stone, C., & Macciò, A. V. 2023, MNRAS, 522, 1208–1227
- Barrientos Acevedo et al. (2023) Barrientos Acevedo, D., van der Wel, A., Baes, M., et al. 2023, MNRAS, 524, 907–922
- Blank et al. (2019) Blank, M., Macciò, A. V., Dutton, A. A., & Obreja, A. 2019, MNRAS, 487, 5476–5489
@@
- Buck et al. (2019) Buck, T., Obreja, A., Macciò, A. V., et al. 2019, MNRAS, 491, 3461
@@
- de los Rios et al. (2023) de los Rios, M., Petač, M., Zaldivar, B., et al. 2023, MNRAS, 525, 6015
@@
- Dutton et al. (2020) Dutton, A. A., Buck, T., Macciò, A. V., et al. 2020, MNRAS, 499, 2648
- El-Badry et al. (2018) El-Badry, K., Bradford, J., Quataert, E., et al. 2018, MNRAS, 477, 1536
@@
- Grand et al. (2024) Grand, R. J. J., Fragkoudi, F., Gómez, F. A., et al. 2024, MNRAS, 532, 1814
@@
- Ho et al. (2024b) Ho, M., Bartlett, D. J., Chartier, N., et al. 2024b, OJAp, 7
- Ho et al. (2019) Ho, M., Rau, M. M., Ntampaka, M., et al. 2019, ApJ, 887, 25
- Hopkins et al. (2018a) Hopkins, P. F., Wetzel, A., Kereš, D., et al. 2018a, MNRAS, 480, 800–863
- Hopkins et al. (2018b) Hopkins, P. F., Wetzel, A., Kereš, D., et al. 2018b, MNRAS, 480, 800–863
@@
- Navarro et al. (1997) Navarro, J. F., Frenk, C. S., & White, S. D. M. 1997, ApJ, 490, 493–508
- Nguyen et al. (2023) Nguyen, T., Mishra-Sharma, S., Williams, R., & Necib, L. 2023, Phys. Rev. D, 107
- Papamakarios et al. (2021) Papamakarios, G., Nalisnick, E., Rezende, D. J., Mohamed, S., & Lakshminarayanan, B. 2021, JMLR, 22, 1
@@
- Paszke et al. (2019) Paszke, A., Gross, S., Massa, F., et al. 2019, PyTorch: An Imperative Style, High-Performance Deep Learning Library
@@
- Sawala et al. (2016) Sawala, T., Frenk, C. S., Fattahi, A., et al. 2016, MNRAS, 457, 1931
- Scarselli et al. (2009) Scarselli, F., Gori, M., Tsoi, A. C., Hagenbuchner, M., & Monfardini, G. 2009, IEEE Trans. Neural Netw., 20, 61
@@
- Shen et al. (2010) Shen, S., Wadsley, J., & Stinson, G. 2010, MNRAS, 407, 1581–1596
@@
- Wang et al. (2015) Wang, L., Dutton, A. A., Stinson, G. S., et al. 2015, MNRAS, 454, 83–94
- Wetzel et al. (2023) Wetzel, A., Hayward, C. C., Sanderson, R. E., et al. 2023, ApJS, 265, 44
@@ Appendix A Bivariate kernel density estimation
Kernel density estimation is a non-parametric technique used to estimate the PDF of a random variable. Bivariate KDE extends this concept to two dimensions, estimating the joint PDF of two random variables.
@@
where …, constitute a sample of size n of the random variable(s) whose PDF we aim to estimate. In the case of bivariate KDE, represents two values , which are joint samples of the two random variables for which we wanted to estimate the PDF. is a kernel function and is a bandwidth matrix whose dimensions match those of (in this case, 2x2).
Effectively, the KDE evaluates the PDF at by adding the contributions of each point , weighted by the kernel function , based on the distance between and , scaled by the bandwidth matrix . We used a 2D Gaussian kernel that follows
where . The bandwidth matrix is calculated by multiplying the covariance matrix of the data by a factor of . Typically, n is the full size of the sample. However, we redefined it as the number of stars inside half the half-light radius ().
@@ Appendix B MAF posterior corner plots
In Fig. 11 we present pairwise contour plots of the posterior samples generated by the CNN+MAF models for the four validation galaxies shown in Fig. 5. These plots illustrate the joint distributions of the inferred values for the enclosed mass at all radii of interest.
Appendix C Predictions for individual radial bins
To better understand the deviations from the true values when predicting the FIRE testing set using our CNN+MAF model trained jointly on the NIHAO and AURIGA simulations, we plot the predicted enclosed mass in each radial bin against the true values in Fig. 12. This figure demonstrates that deviations are generally small in log space. However, we observe a tendency for the model to overpredict the masses of the least massive galaxies in the sample. This behaviour may reflect a bias introduced by the imbalance in our training dataset with respect to enclosed mass (see Fig. 1). Since low-mass galaxies constitute a large fraction of the FIRE testing set, this bias significantly impacts the average model performance when evaluated on this dataset.
We also note a 1.5 dex overprediction for a subset of eight galaxy projections, all corresponding to the same galaxy (which appears 64 times in total, across different projections, in the testing set). Although we ensured that the individual input features (half-light radius, line-of-sight velocity dispersion, and the 98th percentile of line-of-sight velocities) fall within the limits of the training set, it remains possible that some galaxies lie outside the bounds of the training distribution when considering the joint input space. Indeed, we find that projections of this particular galaxy lie near the edges of the input space, suggesting the model is encountering input combinations it has not seen during training. This leads to significant prediction errors. Encouragingly, these outlier cases are associated with large uncertainties in the prediction, allowing us to identify when the inputs are too extreme for the model to produce reliable estimates.
As an alternative approach, we explored a stricter filtering procedure in which we retained only those testing projections that lie within the 99% contours of the training data for every pairwise combination of input features. This method effectively removes the projections associated with large prediction errors, as well as other outliers. However, because these outliers do not significantly affect the overall model performance, we opted not to apply this cut. Instead, we retained all projections to preserve the instructive value of analysing cases where the model fails.
Appendix D Additional posterior calibration checks
@@
We conducted marginal TARP tests on the model trained on NIHAO and AURIGA simulations simultaneously, assessing the calibration of the confidence intervals specifically for each dimension of the 10-d PDF. We show the tests for the validation set and for FIRE galaxy projections in 13. We find similar results to the TARP, demonstrating that the calibration of the posteriors is correct when marginalising over any of its dimensions when validating the model on simulations from the suites used to train it. The posterior distributions produced for FIRE galaxies show similar deviations in their calibration in all dimensions.
@@
After checking the calibration of the errors provided by our CNN+MAF model, we studied their magnitudes. In Fig. 14 we plot the magnitude of the uncertainties estimated by our model for the log10 value of the enclosed mass at different radii, for both the training and the validation set of the model trained on NIHAO and AURIGA simulations simultaneously, where we define the uncertainty as half the range between the 16th and 84th percentile of the enclosed mass at each radii according to the predicted posterior. The median uncertainty peaks for both the validation and training set at 0.125 for the mass enclosed within 0.6Rh, and it decreases monotonically with radius, stabilising at 0.078 for masses within radii equal and larger than 1.8Rh. We remind the reader that posterior distributions generated for FIRE galaxies are not correctly calibrated, and therefore the errors shown for this dataset are not representative of the actual deviation between the predicted and the truth value of the enclosed mass.